OpenStudy (anonymous):

Help with multiple choice please!! 2.8x-2(x+1)=2(3x-1) a.0 b.2 c.no solution d.identity 3. 3(c-2)=2(c-6) a.2 b.-6 c.no solution d.identity 4.0.5(m+4)=3(m-1) a.-2 b.2 c.no solution d.identity

OpenStudy (compassionate):

2.8x-2(x+1)=2(3x-1) 2.8x - 2x - 2 = 6x - 2 Add xs and transpose.

OpenStudy (anonymous):

4x=4x right?

OpenStudy (compassionate):

No.

OpenStudy (anonymous):

oh what is it then?!

OpenStudy (compassionate):

2.8x - 2x - 2 = 6x - 2 0.8x - 2 = 6x - 2 -2 + 2 = 6x - 0.8x

OpenStudy (anonymous):

0=5.2x is what i got

OpenStudy (compassionate):

Correct.

OpenStudy (compassionate):

Therefore, there is?

OpenStudy (anonymous):

0, so no solution?

OpenStudy (compassionate):

Yes.

OpenStudy (anonymous):

okay thank you

OpenStudy (compassionate):

What about: 3(c-2)=2(c-6)

OpenStudy (anonymous):

not sure althoughi think its identity

OpenStudy (compassionate):

Yes.

OpenStudy (anonymous):

okay ill figure the last one out thanks for all the help(: